# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4.3 Coupled Toda-DST system
$L_{\text{Toda-DST}}(\lambda)=L_{\text{Toda}}(\lambda)+\beta L_{\text{DST}}(\lambda)\,,$ (4.46)
where $L_{\text{Toda}}(\lambda)$ is the Lax matrix of the periodic Toda chain in (4.9), $L_{\text{DST}}(\lambda)$ is the DST Lax matrix (4.31), and the parameter $\beta$ is a real-valued scalar parameter dictating the strength of coupling between the two hierarchies. Naturally, the Lax matrix $L_{\text{Toda-DST}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix:
$L_{\text{Toda-DST}}(\lambda)=\frac{J_{0}^{(0)}+\beta K_{0}^{(0)}}{\lambda}+\frac{J_{0}^{(1)}}{\lambda^{2}}+\frac{\beta}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}+J_{\infty}+\beta K_{\infty}$ (4.47)
@@
Orbit realisation:
Set $D=\{0,1,\infty\}$. Since we already have a parameterisation of the Lax matrices $L_{\text{Toda}}(\lambda)$ and $L_{\text{DST}}(\lambda)$ as orbit elements, we only need to check that the action of a generic group element $\varphi_{+}=\left(\varphi_{0+},\varphi_{1+},\varphi_{\infty+}\right)\in\widetilde{G}_{D+}^{(0)}$ defined by (3.47) on the non-dynamical element
$\Lambda_{\text{Toda-DST}}(\lambda)=\Lambda_{\text{Toda}}(\lambda)+\beta\Lambda_{\text{DST}}(\lambda)\in\Omega_{D}^{\Gamma}$ (4.48)
@@
$\Lambda_{\text{Toda}}(\lambda)=\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\Lambda_{\infty}\,,\quad\Lambda_{\text{DST}}(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{1}}{\lambda-\omega^{k}\zeta_{1}}+\Lambda_{\infty}$ (4.49)
@@
$\begin{split}&J_{0}^{(0)}+\beta K_{0}^{(0)}=[\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},J_{0}^{(1)}]+\beta\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1}\,,\\ &J_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1}\,,\\ &\beta K_{1}=\beta\phi_{1}\Lambda_{1}\phi_{1}^{-1}\,,\quad\text{and}\\ &J_{\infty}+\beta K_{\infty}=\Lambda_{\infty}+\beta\Lambda_{\infty}\,,\end{split}$ (4.50)
as desired. This gives us a parameterisation of the Lax matrix $L_{\text{Toda-DST}}(\lambda)$ in (4.46) as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{\text{Toda}}\times{\cal O}_{\Lambda}^{\text{DST}}$.
Lax equations:
As earlier, we will choose invariant functions $H_{p,r}$ on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ as defined in (3.70), for $r\in\{0,1\}$. The Lax equations for the DST-Toda model with respect to the elementary times $t_{p}^{r}$ are given by
$\partial_{t_{p}^{r}}\iota_{\lambda}L_{\text{Toda-DST}}=[R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\iota_{\lambda}L_{\text{Toda-DST}}]\,.$ (4.51)
@@
$M_{1,r}=\begin{dcases}-\dfrac{J_{0}^{(1)}}{\lambda}&\text{for}\,\,r=0\\ -\dfrac{\beta}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{1}\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}&\text{for}\,\,r=1\,.\end{dcases}$
@@
Taking residues on both sides in (4.52) gives the equations of motion for $p_{i},q_{i},X_{i},x_{i}$, for $i=1,\ldots,T$. Like in the cases of the periodic Toda chain and the DST model, to get the equations of motion for $q_{i},X_{i},x_{i}$, one can “undo” the dressing in the corresponding equations. Explicitly, one gets the following equations for $p_{i},q_{i}$, for $i=1,\ldots,T$:
$\begin{cases}\partial_{t_{1}^{0}}p_{i}=(1+\beta)(e^{q_{i}-q_{i+1}}-e^{q_{i-1}-q_{i}})+\dfrac{\beta}{\zeta_{1}}(e^{q_{i-1}-q_{i}}x_{i-1}X_{i}-e^{q_{i}-q_{i+1}}x_{i}X_{i+1})\,,\\[6.45831pt] \partial_{t_{1}^{0}}q_{i}=-p_{i}-\beta c_{i}\,,\\[6.45831pt] \partial_{t_{1}^{1}}p_{i}=\dfrac{\beta}{\zeta_{1}}(e^{q_{i}-q_{i+1}}x_{i}X_{i+1}-e^{q_{i-1}-q_{i}}x_{i-1}X_{i})\,,\\[6.45831pt] \partial_{t_{1}^{1}}q_{i}=-\beta x_{i}X_{i}\,,\end{cases}$ (4.53)
and the following equations for $X_{i},x_{i}$, for $i=1,\ldots,T$:
$\begin{cases}\partial_{t_{1}^{0}}X_{i}=\dfrac{1}{\zeta_{1}}e^{q_{i}-q_{i+1}}X_{i+1}\,,\\[6.45831pt] \partial_{t_{1}^{0}}x_{i}=-\dfrac{1}{\zeta_{1}}e^{q_{i-1}-q_{i}}x_{i-1}\,,\\[6.45831pt] \partial_{t_{1}^{1}}X_{i}=-p_{i}X_{i}-\beta c_{i}X_{i}-\dfrac{1}{\zeta_{1}}e^{q_{i}-q_{i+1}}X_{i+1}-\dfrac{\beta}{T}\sum\limits_{j=1}^{T}\sum\limits_{k=0}^{T-1}\omega^{k(j-i)}x_{j}X_{j}X_{i}-(1+\beta)\zeta_{1}X_{i-1}\,,\\[6.45831pt] \partial_{t_{1}^{1}}x_{i}=p_{i}x_{i}+\beta c_{i}x_{i}+\dfrac{1}{\zeta_{1}}e^{q_{i-1}-q_{i}}x_{i-1}+\dfrac{\beta}{T}\sum\limits_{j=1}^{T}\sum\limits_{k=0}^{T-1}\omega^{k(j-i)}X_{j}x_{j}x_{i}+(1+\beta)\zeta_{1}x_{i+1}\,.\end{cases}$ (4.54)
@@
Setting $\beta=0$ in the above equations produces the equations of motion in (4.21) for the periodic Toda chain. In the limit $\beta\rightarrow\infty$, these reduce to the equations of motion we obtained for the DST model in (4.42). To see this, note that the Lax matrix $L_{\text{DST}}$ comes as $\beta L_{\text{DST}}$ in the coupled system. Therefore, in the limit $\beta\rightarrow\infty$, the time flow $t_{p}^{r}$ is rescaled such that $\partial_{t_{p}^{r}}$ rescales to $\beta^{p}\partial_{t_{p}^{r}}$.
Lagrangian description:
@@
Using Theorem 3.6, we can now write a Lagrangian multiform for the Toda-DST hierarchy:
$\mathscr{L}_{\text{Toda-DST}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\mathrm{d}t_{p}^{0}+\sum_{p=1}^{M}\mathscr{L}_{p,1}\mathrm{d}t_{p}^{1}$ (4.55)
where
$\mathscr{L}_{p,r}=\operatorname{Tr}(J_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1})+\beta\operatorname{Tr}(K_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1})$
@@
$\qquad\quad+\beta\operatorname{Tr}(K_{1}\partial_{t_{p}^{r}}\phi_{1}\phi_{1}^{-1})-H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\quad r\in\{0,1\}\,.$ (4.56)
@@
where the potential term $H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}})$ is given by (3.70) for $r\in\{0,1\}$. Notice that $K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}$. So, the second term on the right-hand side is, in fact, a total derivative and will not enter the Euler-Lagrange equations. So, we will simply drop it. In terms of the canonical coordinates, we then have
@@
$\mathscr{L}_{p,r}=-\sum_{i=1}^{T}p_{i}\partial_{t_{p}^{r}}q_{i}+\beta\sum_{i=1}^{T}X_{i}\partial_{t_{p}^{r}}x_{i}-H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\quad r\in\{0,1\}\,.$ (4.57)
@@
The decoupled limits of the periodic Toda and the DST hierarchies are easily recovered by setting $\beta=0$ and taking the limit $\beta\rightarrow\infty$ in (4.57) respectively.
@@
$\begin{split}&H_{1,0}=\frac{1}{2}\sum_{i=1}^{T}p_{i}^{2}+\beta\sum_{i=1}^{T}c_{i}p_{i}+\frac{\beta^{2}}{2}\sum_{i=1}^{T}c_{i}^{2}+(1+\beta)\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}-\frac{\beta}{\zeta_{1}}\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}x_{i}X_{i+1}\,,\\ &H_{1,1}=\frac{\beta^{2}}{2T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\omega^{k(j-i)}x_{i}x_{j}X_{i}X_{j}+\beta\sum_{i=1}^{T}p_{i}x_{i}X_{i}+\beta^{2}\sum_{i=1}^{T}c_{i}x_{i}X_{i}\\ &\qquad\quad+\frac{\beta}{\zeta_{1}}\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}x_{i}X_{i+1}+(\beta+\beta^{2})\zeta_{1}\sum_{i=1}^{T}x_{i+1}X_{i}\,.\end{split}$ (4.58)
The variation $\delta\mathscr{L}_{1,r}$ gives the Euler-Lagrange equations for the coupled system with respect to the time flow $t_{1}^{r}$, for each $r\in\{0,1,\infty\}$. It can be checked that these are exactly the equations we obtained in (4.53)-(4.54).
@@ 5 Conclusion
From the point of view of Lagrangian multiforms, our main result is the construction of a Lagrangian multiform for the class of cyclotomic Gaudin models which includes, as a particular example, the emblematic periodic Toda chain. The main technical aspect of this result lies in the pinning down of the necessary algebraic setup (in Section 3) associated with the classical $r$-matrix of the cyclotomic Gaudin model which is both non-skew-symmetric and depends on spectral parameters. This is the first time that a Lagrangian multiform has been constructed for an $r$-matrix with both of these properties, as previous general constructions were either only presented for skew-symmetric ones [CSV, CDS] or for non-skew-symmetric $r$-matrices but without spectral parameter [CDS].
The non-skew-symmetric $r$-matrix of the cyclotomic Gaudin model arises from a decomposition of a direct sum of (twisted) loop algebras into a pair of complementary subalgebras, which controls the corresponding Lie dialgebra structure entering its Lagrangian multiform. While finishing this work, the preprint [AMS] appeared on the construction of solutions of the general classical Yang–Baxter equation based on more general decompositions of such Lie algebras and to which one can associate generalised notions of Gaudin models. In particular, it appears that the cyclotomic setup considered here is a special case of their construction. It would be interesting to investigate if our Lagrangian multiform construction can be extended to the more general non-skew-symmetric $r$-matrices and associated generalised Gaudin models of [AMS].

### v2

@@ 4.3 Coupled Toda-DST system
$L_{\text{Toda-DST}}(\lambda)=L_{\text{Toda}}(\lambda)+\beta L_{{\rm DST}}(\lambda),$ (4.14)
where $L_{\text{Toda}}(\lambda)$ is the Lax matrix of the periodic Toda chain in (4.5), $L_{{\rm DST}}(\lambda)$ is the DST Lax matrix (4.10), and the parameter $\beta$ is a real-valued scalar parameter dictating the strength of coupling between the two hierarchies. Naturally, the Lax matrix $L_{\text{Toda-DST}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix:
$L_{\text{Toda-DST}}(\lambda)=\frac{J_{0}^{(0)}+\beta K_{0}^{(0)}}{\lambda}+\frac{J_{0}^{(1)}}{\lambda^{2}}+\frac{\beta}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}+J_{\infty}+\beta K_{\infty}$
@@
Orbit realisation. Set $D=\{0,1,\infty\}$. Since we already have a parameterisation of the Lax matrices $L_{\text{Toda}}(\lambda)$ and $L_{{\rm DST}}(\lambda)$ as orbit elements, we only need to check that the action of a generic group element $\varphi_{+}=(\varphi_{0+},\varphi_{1+},\varphi_{\infty+})\in\widetilde{G}_{D+}^{(0)}$ defined by (3.24) on the non-dynamical element
$\Lambda_{\text{Toda-DST}}(\lambda)=\Lambda_{\text{Toda}}(\lambda)+\beta\Lambda_{{\rm DST}}(\lambda)\in\Omega_{D^{\prime}}^{\Gamma},$
@@
$\Lambda_{\text{Toda}}(\lambda)=\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\Lambda_{\infty},\qquad\Lambda_{{\rm DST}}(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{1}}{\lambda-\omega^{k}\zeta_{1}}+\Lambda_{\infty}$
@@
$J_{0}^{(0)}+\beta K_{0}^{(0)}=\bigl{[}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},J_{0}^{(1)}\bigr{]}+\beta\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1},$
$J_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1},$
$\beta K_{1}=\beta\phi_{1}\Lambda_{1}\phi_{1}^{-1},\qquad\text{and}$
$J_{\infty}+\beta K_{\infty}=\Lambda_{\infty}+\beta\Lambda_{\infty},$
as desired. This gives us a parameterisation of the Lax matrix $L_{\text{Toda-DST}}(\lambda)$ in (4.14) as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{\text{Toda}}\times{\cal O}_{\Lambda}^{{\rm DST}}$.
Lax equations. As earlier, we will choose invariant functions $H_{p,r}$ on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ as defined in (3.36), for $r\in\{0,1\}$. The Lax equations for the DST-Toda model with respect to the elementary times $t_{p}^{r}$ are given by
$\partial_{t_{p}^{r}}\iota_{\lambda}L_{\text{Toda-DST}}=[R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\iota_{\lambda}L_{\text{Toda-DST}}].$
@@
$M_{1,r}=\begin{dcases}-\dfrac{J_{0}^{(1)}}{\lambda}&\text{for}\ r=0,\\ -\dfrac{\beta}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{1}\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}&\text{for}\ r=1.\end{dcases}$
@@
Taking residues on both sides in (4.15) gives the equations of motion for $p_{i}$, $q_{i}$, $X_{i}$, $x_{i}$, for ${i=1,\dots,T}$. Like in the cases of the periodic Toda chain and the DST model, to get the equations of motion for $q_{i}$, $X_{i}$, $x_{i}$, one can “undo” the dressing in the corresponding equations. Explicitly, one gets the following equations for $p_{i}$, $q_{i}$ for $i=1,\dots,T$:
$\partial_{t_{1}^{0}}p_{i}=(1+\beta)\bigl{(}{\rm e}^{q_{i}-q_{i+1}}-{\rm e}^{q_{i-1}-q_{i}}\bigr{)}+\dfrac{\beta}{\zeta_{1}}\bigl{(}{\rm e}^{q_{i-1}-q_{i}}x_{i-1}X_{i}-{\rm e}^{q_{i}-q_{i+1}}x_{i}X_{i+1}\bigr{)},$
$\partial_{t_{1}^{0}}q_{i}=-p_{i}-\beta c_{i},$
$\partial_{t_{1}^{1}}p_{i}=\dfrac{\beta}{\zeta_{1}}\bigl{(}{\rm e}^{q_{i}-q_{i+1}}x_{i}X_{i+1}-{\rm e}^{q_{i-1}-q_{i}}x_{i-1}X_{i}\bigr{)},$
$\partial_{t_{1}^{1}}q_{i}=-\beta x_{i}X_{i},$ (4.16)
and the following equations for $X_{i}$, $x_{i}$, for $i=1,\dots,T$:
$\partial_{t_{1}^{0}}X_{i}=\dfrac{1}{\zeta_{1}}{\rm e}^{q_{i}-q_{i+1}}X_{i+1},$
$\partial_{t_{1}^{0}}x_{i}=-\dfrac{1}{\zeta_{1}}{\rm e}^{q_{i-1}-q_{i}}x_{i-1},$
$\partial_{t_{1}^{1}}X_{i}=-p_{i}X_{i}-\beta c_{i}X_{i}-\dfrac{1}{\zeta_{1}}{\rm e}^{q_{i}-q_{i+1}}X_{i+1}-\dfrac{\beta}{T}\sum\limits_{j=1}^{T}\sum\limits_{k=0}^{T-1}\omega^{k(j-i)}x_{j}X_{j}X_{i}-(1+\beta)\zeta_{1}X_{i-1},$
$\partial_{t_{1}^{1}}x_{i}=p_{i}x_{i}+\beta c_{i}x_{i}+\dfrac{1}{\zeta_{1}}{\rm e}^{q_{i-1}-q_{i}}x_{i-1}+\dfrac{\beta}{T}\sum\limits_{j=1}^{T}\sum\limits_{k=0}^{T-1}\omega^{k(j-i)}X_{j}x_{j}x_{i}+(1+\beta)\zeta_{1}x_{i+1}.$ (4.17)
@@
Setting $\beta=0$ in the above equations produces the equations of motion in (4.8) for the periodic Toda chain. In the limit $\beta\rightarrow\infty$, these reduce to the equations of motion we obtained for the DST model in (4.13). To see this, note that the Lax matrix $L_{{\rm DST}}$ comes as $\beta L_{{\rm DST}}$ in the coupled system. Therefore, in the limit $\beta\rightarrow\infty$, the time flow $t_{p}^{r}$ is rescaled such that $\partial_{t_{p}^{r}}$ rescales to $\beta^{p}\partial_{t_{p}^{r}}$.
@@
Lagrangian description. Using Theorem 3.6, we can now write a Lagrangian multiform for the Toda-DST hierarchy:
$\mathscr{L}_{\text{Toda-DST}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\mathrm{d}t_{p}^{0}+\sum_{p=1}^{M}\mathscr{L}_{p,1}\mathrm{d}t_{p}^{1}$
with
$\mathscr{L}_{p,r}$ ${}=\operatorname{Tr}\bigl{(}J_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\bigr{)}+\beta\operatorname{Tr}\bigl{(}K_{0}^{(0)}\partial_{t_{p}^{r}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}\bigr{)}$
@@
$\quad{}+\beta\operatorname{Tr}\bigl{(}K_{1}\partial_{t_{p}^{r}}\phi_{1}\phi_{1}^{-1}\bigr{)}-H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\qquad r\in\{0,1\},$
@@
where the potential term $H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}})$ is given by (3.36) for $r\in\{0,1\}$. Notice that $K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}$. So, the second term on the right-hand side is, in fact, a total derivative and will not enter the Euler–Lagrange equations. So, we will simply drop it. In terms of the canonical coordinates, we then have
@@
$\mathscr{L}_{p,r}=-\sum_{i=1}^{T}p_{i}\partial_{t_{p}^{r}}q_{i}+\beta\sum_{i=1}^{T}X_{i}\partial_{t_{p}^{r}}x_{i}-H_{p,r}(\iota_{\lambda}L_{\text{Toda-DST}}),\qquad r\in\{0,1\}.$ (4.18)
@@
The decoupled limits of the periodic Toda and the DST hierarchies are easily recovered by setting $\beta=0$ and taking the limit $\beta\rightarrow\infty$ in (4.18), respectively.
@@
$H_{1,0}=\frac{1}{2}\sum_{i=1}^{T}p_{i}^{2}+\beta\sum_{i=1}^{T}c_{i}p_{i}+\frac{\beta^{2}}{2}\sum_{i=1}^{T}c_{i}^{2}+(1+\beta)\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}-\frac{\beta}{\zeta_{1}}\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}x_{i}X_{i+1},$
$H_{1,1}=\frac{\beta^{2}}{2T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\omega^{k(j-i)}x_{i}x_{j}X_{i}X_{j}+\beta\sum_{i=1}^{T}p_{i}x_{i}X_{i}+\beta^{2}\sum_{i=1}^{T}c_{i}x_{i}X_{i}$
${}+\frac{\beta}{\zeta_{1}}\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}x_{i}X_{i+1}+\bigl{(}\beta+\beta^{2}\bigr{)}\zeta_{1}\sum_{i=1}^{T}x_{i+1}X_{i}.$
The variation $\delta\mathscr{L}_{1,r}$ gives the Euler–Lagrange equations for the coupled system with respect to the time flow $t_{1}^{r}$, for each $r\in\{0,1\}$. It can be checked that these are exactly the equations we obtained in (4.16) and (4.17).
@@ 5 Conclusion
From the point of view of Lagrangian multiforms, our main result is the construction of a Lagrangian multiform for the class of cyclotomic Gaudin models which includes, as a particular example, the emblematic periodic Toda chain. The main technical aspect of this result lies in the pinning down of the necessary algebraic setup (in Section 3) associated with the classical $r$-matrix of the cyclotomic Gaudin model which is both non-skew-symmetric and depends on spectral parameters. This is the first time that a Lagrangian multiform has been constructed for an $r$-matrix with both of these properties, as previous general constructions were either only presented for skew-symmetric ones [11, 15] or for non-skew-symmetric $r$-matrices but without spectral parameter [11].
The non-skew-symmetric $r$-matrix of the cyclotomic Gaudin model arises from a decomposition of a direct sum of (twisted) loop algebras into a pair of complementary subalgebras, which controls the corresponding Lie dialgebra structure entering its Lagrangian multiform. While finishing this work, the preprint [1] appeared on the construction of solutions of the general classical Yang–Baxter equation based on more general decompositions of such Lie algebras and to which one can associate generalised notions of Gaudin models. In particular, it appears that the cyclotomic setup considered here is a special case of their construction. It would be interesting to investigate if our Lagrangian multiform construction can be extended to the more general non-skew-symmetric $r$-matrices and associated generalised Gaudin models of [1].